# Brought a Gun to a Knife fight: Modern VFM Baselines Outgun Specialized Detectors on In-the-Wild AI Image Detection

### v2

@@ 1 Introduction
As a result, a central challenge for AIGI detection is building models with a strong generalization ability, allowing them to effectively identify and verify images generated by a variety of unknown methods. To address this, researchers have developed various forensics-specialized detectors, such as incorporating additional augmented samples (e.g., full or partial image reconstructions) [1, 2, 3] or designing improved training strategies [4]. These approaches have already pushed the performance on recent AIGI benchmarks to impressive levels. Consequently, the research focus has begun to shift toward the more challenging in-the-wild datasets. However, such real-world scenarios remain highly challenging, as most forensic methods fail to reach satisfactory performance, particularly on the Chameleon [5], as illustrated in Fig. 1.
@@
To investigate this issue, we conduct two analyses. We first test all VFMs on a verifiably unseen dataset composed of post-cutoff synthetic media and private, personally captured photographs. As hypothesized, the performance of most modern VFMs drops significantly, revealing their reliance on prior data exposure. Building on this, our second analysis probes the underlying mechanism for Vision-Language Models. We find they have learned to align synthetic images with forgery-related concepts (e.g., ’AI-generated’), providing a semantic explanation for their effectiveness. This ’built-in’ knowledge, learned from web-scale data, allows these general-purpose models to decisively outgun specialist detectors explicitly designed for the forensic task.
@@ 3.1 Testing for Data Exposure with Verifiably Unseen Data
Method Chameleon SocialRF Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) modern design 0.628 portrait 0.345 forged 0.332 urban 0.253 Siglip (2023.3.27) unaltered 0.318 original 0.212 edited 0.232 original 0.222 Siglip2 (2025.2.21) genuine 0.385 urban 0.250 portrait 0.202 vintage 0.191 Meta CLIP (2023.9.28) AI generated 0.678 deepfake 0.091 AI generated 0.902 original 0.024 Meta CLIP-2 (2025.7.29) AI generated 0.828 deepfake 0.064 AI generated 0.924 deepfake 0.038 PE (2025.4.17) AI generated 0.861 deepfake 0.021 AI generated 0.943 deepfake 0.031 Method CommunityAI WebAIG-25 Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) portrait 0.346 nature 0.326 urban 0.332 midjourney_images 0.260 Siglip (2023.3.27) real 0.281 AIGIBench 0.283 fake 0.278 original 0.246 Siglip2 (2025.2.21) urban 0.209 portrait 0.201 urban 0.212 portrait 0.194 Meta CLIP (2023.9.28) AI generated 0.726 deepfake 0.086 midjourney_images 0.604 AI generated 0.284 Meta CLIP-2 (2025.7.29) AI generated 0.858 deepfake 0.041 midjourney_images 0.621 AI generated 0.308 PE (2025.4.17) AI generated 0.878 CommunityAI 0.029 midjourney_images 0.722 AI generated 0.218

### v3

@@ 1 Introduction
As a result, a central challenge for AIGI detection is building models with a strong generalization ability, allowing them to effectively identify and verify images generated by a variety of unknown methods. To address this, researchers have developed various forensics-specialized detectors, such as incorporating additional augmented samples (e.g., full or partial image reconstructions) [1, 2, 3] or designing improved training strategies [4, 5, 6]. These approaches have already pushed the performance on recent AIGI benchmarks to impressive levels. Consequently, the research focus has begun to shift toward the more challenging in-the-wild datasets. However, such real-world scenarios remain highly challenging, as most forensic methods fail to reach satisfactory performance, particularly on the Chameleon [7], as illustrated in Fig. 1.
@@
To investigate this issue, we conduct two analyses. We first test all VFMs on a verifiably unseen dataset composed of post-cutoff synthetic media and private, personally captured photographs. As hypothesized, the performance of most modern VFMs drops significantly, revealing their reliance on prior data exposure. Building on this, our second analysis probes the underlying mechanism for Vision-Language Models. We find they have learned to align synthetic images with forgery-related concepts (e.g., ‘AI-generated’), providing a semantic explanation for their effectiveness. This ’built-in’ knowledge, learned from web-scale data, allows these general-purpose models to decisively outgun specialist detectors explicitly designed for the forensic task.
@@ 3.1 Testing for Data Exposure with Verifiably Unseen Data
Method Chameleon SocialRF Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) modern design 0.628 portrait 0.345 forged 0.332 urban 0.253 Siglip (2023.3.27) unaltered 0.318 original 0.212 edited 0.232 original 0.222 Siglip2 (2025.2.21) genuine 0.385 urban 0.250 portrait 0.202 vintage 0.191 Meta CLIP (2023.9.28) AI generated 0.678 deepfake 0.091 AI generated 0.902 original 0.024 Meta CLIP-2 (2025.7.29) AI generated 0.828 deepfake 0.064 AI generated 0.924 deepfake 0.038 PE (2025.4.17) AI generated 0.861 deepfake 0.021 AI generated 0.943 deepfake 0.031 Method CommunityAI WebAIG-25 Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) portrait 0.346 nature 0.326 urban 0.332 midjourney_images 0.260 Siglip (2023.3.27) AIGIBench 0.283 real 0.281 fake 0.278 original 0.246 Siglip2 (2025.2.21) urban 0.209 portrait 0.201 urban 0.212 portrait 0.194 Meta CLIP (2023.9.28) AI generated 0.726 deepfake 0.086 midjourney_images 0.604 AI generated 0.284 Meta CLIP-2 (2025.7.29) AI generated 0.858 deepfake 0.041 midjourney_images 0.621 AI generated 0.308 PE (2025.4.17) AI generated 0.878 CommunityAI 0.029 midjourney_images 0.722 AI generated 0.218